If the ratio of 7a to 8b is 9 to 10, what is the ratio of 10a to 7b?
Answer by Alan3354(69443)

If the ratio of 7a to 8b is 9 to 10, what is the ratio of 10a to 7b?
---------
7a/8b = 9/10
70a = 72b
----
10a = 7.2b
10a/7b = 7.2b/7b = 7.2/7
10a/7b = 36/35